# Deep Learning for Protein-Ligand Docking: Are We There Yet?

### v2

@@ 3.1 Preprocessed datasets
To perform apo docking with this dataset, we used ESMFold (Lin et al., 2023) to predict the complex structure of each of its proteins, where 5 of these 85 complexes were excluded from the effective benchmarking set due to being too large for structure prediction on an 80GB NVIDIA A100 GPU. For the remaining 80 complexes, we then optimally aligned their predicted protein structures to the corresponding ground-truth (holo) protein-ligand structures using the PLI-weighted RMSD alignment algorithm originally proposed by Corso et al. (2022).
@@
Likewise to Astex Diverse, for the PoseBusters Benchmark set, we used ESMFold to predict the apo complex structures of each of its proteins. After filtering out 28 complexes for which the corresponding protein structure could not be predicted on an 80GB A100 GPU, we RMSD-aligned the remaining 280 predicted protein structures while optimally weighting each complex’s protein-ligand interface in the alignment. For the DockGen dataset, we refer readers to Appendix F.1.
@@
Sequence identity overlap. Note that for all four of the test datasets described above and listed in Table 1, we do not perform an analysis of the sequence identity overlap between these test datasets’ proteins and those of e.g., the PDBBind (time-split) training dataset, as (1) not all DL-based docking methods use PDBBind as their respective training datasets and (2) leaving the test complexes unfiltered according to sequence identity should, in principle, reflect many real-world use cases of these methods in which several (new) protein targets they are presented with may or may not be similar to what the methods have "seen" during training. Nevertheless, for an investigation of the sequence identity overlap between e.g., the PoseBusters Benchmark set and PDBBind, we refer interested readers to Buttenschoen et al. (2024). Furthermore, in Appendix E, we analyze the different types and frequencies of protein-ligand interactions natively found within the Astex Diverse, PoseBusters Benchmark, DockGen, and CASP15 datasets, respectively, to quantify the diversity of the (predicted) interactions each dataset can be used to evaluate.
@@ 4 Methods and experimental setup
[enumerate] \1 Formats for conventional methods are as follows: \2 Template-based methods such as TULIP are provided with an apo (predicted) protein structure and (fragment) ligand SMILES strings and are tasked with retrieving (PDB template (Bank, 1971)) ligand conformations residing in the same coordinate system as the given (predicted) protein structure following optimal molecular and structural alignment (Hu et al., 2018) with corresponding RDKit conformers of the input (query) ligand SMILES strings, where structural similarity with the query ligands is used to rank-order the selected (PDB template) conformations. \2 Molecular docking tools such as AutoDock Vina, which require specification of protein binding sites, are provided with not only a predicted protein structure but also the centroid coordinates of each (DiffDock-L-)predicted protein-ligand binding site residue. Such binding site residues are classified using a 4 Å protein-ligand heavy atom interaction threshold and using a 25 Å ligand-ligand heavy atom interaction threshold to define a "group" of ligands belonging to the same binding site and therefore residing in the same 25 Å3-sized binding site input voxel for AutoDock Vina.
@@
Formats for generative methods are as follows: \2 DiffDock-L is provided with a predicted protein structure and (fragment) ligand SMILES strings and is then tasked with producing (multiple rank-ordered) ligand conformations (for each fragment) for the given protein. Note that DiffDock-L does not natively support multi-ligand SMILES string inputs, so in this work, we propose a modified inference procedure for DiffDock-L which autoregressively presents each (fragment) ligand SMILES string to the model while providing the same predicted protein structure to the model in each inference iteration (reporting for each complex the average confidence score over all iterations). Notably, as an inference-time modification, this sampling formulation permits multi-ligand sampling yet cannot model multi-ligand interactions directly and therefore often produces ligand-ligand steric clashes. \2 As a single-ligand generative docking method, DynamicBind adopts the same input and output formats as DiffDock-L with the following exceptions: (1) the predicted input protein structure can be modified in response to (fragment) ligand docking; (2) the autoregressive inference procedure we adapted from that of DiffDock-L now provides DynamicBind with its own most recently generated protein structure in each (fragment) ligand inference iteration, thereby providing the model with partial multi-ligand interaction context; and (3) iteration-averaged confidence scores and predicted affinities are reported for each complex. Nonetheless, for both DiffDock-L and DynamicBind, such modified inference procedures highlight the importance in future work of retraining such generative methods directly on multi-ligand complexes to address such inference-time compromises. \2 Lastly, as a natively multi-ligand structure generation model pretrained using various 3D molecular and protein data sources, NeuralPLexer receives as its inputs a (multi-chain) protein sequence, a predicted protein (template) structure, as well as (fragment) ligand SMILES strings. The method is then tasked with producing (multiple rank-ordered) protein-ligand structure conformations for each input complex, using the method’s average predicted per-ligand heavy atom lDDT score (Mariani et al., 2013) for rank-ordering.
@@
Multi-ligand evaluation. Following CASP15’s official scoring procedure for protein-ligand complexes (Robin et al., 2023), for multi-ligand targets, we report each method’s percentage of "correct" (binding site-superimposed) ligand conformations (RMSD $\leq$ 2 Å) as well as violin plots of the RMSD and PLI-specific lDDT scores of its protein-ligand conformations across all (fragment) ligands within the benchmark’s multi-ligand complexes (see Appendix F for these plots). Notably, this final metric, referred to lDDT-PLI, allows one to evaluate specifically how well each method can model protein-ligand structural interfaces. In the remainder of this work, we will discuss our benchmark’s results and their implications for the development of future complex structure generation methods.
@@ 5 Results and discussions
In this section, we present PoseBench’s results for single and multi-ligand protein-ligand structure generation and discuss their implications for future work. Note that across all the experiments, for generative methods (or methods that use generative inputs to make their predictions), we report their performance metrics in terms of the mean and standard deviation across three independent runs of the method to gain insights into its inter-run stability and consistency.
@@ 5.1 Generalization to new binding pockets implies single-ligand docking performance
We begin our investigations by evaluating the performance of each baseline method for single-ligand docking using the Astex Diverse and PoseBusters Benchmark datasets. Notably, for results on the PoseBusters Benchmark dataset, we perform an additional analysis where we apply post-prediction (fixed-protein) relaxation to each method’s generated ligand conformations using molecular dynamics simulations (Eastman and Pande, 2010), as originally proposed by Buttenschoen et al. (2024). As shown in Figures 2 and 3, DiffDock-L achieves the best overall performance across the two datasets both with and without applying relaxation to its generated structures. Closely behind in performance for the PoseBusters Benchmark dataset are DynamicBind and RoseTTAFold-All-Atom following structural relaxation. Interestingly, without relaxation, AutoDock Vina combined with DiffDock-L’s predicted binding pockets achieves the second-best performance on the PoseBusters Benchmark dataset, which suggests that DiffDock-L is currently the only single-ligand deep learning method that presents a better intrinsic understanding of biomolecular physics for docking than conventional modeling tools. For interested readers, in Appendix F.2, we report e.g., pocket-only PoseBusters Benchmark experiments and RMSD violin plots for both the Astex Diverse and PoseBusters Benchmark datasets.
@@ 5.2 Molecule pretraining implies multi-ligand docking performance
Another interesting observation is that TULIP (following structural relaxation) outperforms the docking success rates of single-ligand deep learning docking methods such as DiffDock-L and DynamicBind in the context of multi-ligand docking, which suggests room for future improvement in the multi-ligand modeling capabilities of these recent deep learning baselines. To further inspect each method’s understanding of biomolecular physics for docking, in Figure 5 we report each method’s percentage of predicted complexes (whether correct or not) for which all ligand conformations in the complex are jointly considered valid according to the PoseBusters software suite (i.e., PB-Valid). In short, in the context of multi-ligands, we find that AutoDock Vina followed by DiffDock-L are tied in terms of their PoseBusters validity rates following structural relaxation, with our ensembling consensus baseline (i.e., Ensemble (Con)) as well as DynamicBind shortly behind. As an addendum, we note that NeuralPLexer’s (DynamicBind’s) predictions seem to be frequently selected by Ensemble (Con) for single (multi)-ligand targets, which suggests that NeuralPLexer (DynamicBind) produces the most consistent (i.e., similar) ligand poses for a given single (multi)-ligand protein complex. For interested readers, in Appendix F.3, we report additional results e.g., in terms of lDDT-PLI and RMSD violin plots for both the total available CASP15 targets as well as the publicly available ones.
@@ Conclusions
In this work, we introduced PoseBench, the first deep learning benchmark for practical protein-ligand docking. Experimental results with PoseBench suggest the importance of developing new multi-ligand structure generation methods for enhanced generalization in future work. Key limitations of this study include its reliance on the accuracy of its predicted protein structures, its limited number of multi-ligand prediction targets available for benchmarking, and its inclusion of only a subset of all available protein-ligand docking baselines to focus on the most recent deep learning algorithms designed specifically for docking. In future work, we aim to expand not only the number of baseline methods but also the number of available multi-ligand targets while maintaining a diverse composition of heterogeneous (ionic) complexes. As a publicly available resource, PoseBench is flexible to accommodate new datasets and methods for protein-ligand structure generation.
@@ Appendix C Compute resources
Appendix D Documentation for datasets
@@
D.1 Astex Diverse Set - Single-Ligand Docking (Difficulty: Easy)
@@
D.2 PoseBusters Benchmark Set - Single-Ligand Docking (Difficulty: Intermediate)
@@
D.3 DockGen Set - Single-Ligand Docking (Difficulty: Challenging)
@@
D.4 CASP15 Set - Multi-Ligand Docking (Difficulty: Challenging)
@@
Appendix E Analysis of protein-ligand interactions
@@
Appendix F Additional results
@@ F.1 DockGen results
Benchmark results. Figures 8 and 9 show that DiffDock-L, NeuralPLexer, and Ensemble (Con) provide the best pocket generalization capabilities compared to all other baseline methods. In particular, Ensemble (Con) seems to provide the most stable performance in this setting. However, we note that none of the baseline methods generate any PB-valid ligand conformations, suggesting that such "correct" poses are approximately accurate yet physically implausible in certain measurable ways.
@@
F.2 Expanded Astex & PoseBusters results
F.2.1 Pocket-only PoseBusters results

### v3

@@ 3.1 Preprocessed datasets
To perform apo docking with this dataset, we used ESMFold (Lin et al., 2023) to predict the complex structure of each of its proteins, where 5 of these 85 complexes were excluded from the effective benchmarking set due to being too large for structure prediction on an 80GB NVIDIA A100 GPU. For the remaining 80 complexes, we then optimally aligned their predicted protein structures to the corresponding ground-truth (holo) protein-ligand structures using the PLI-weighted root mean square deviation (RMSD) alignment algorithm originally proposed by Corso et al. (2022).
@@
Likewise to Astex Diverse, for the PoseBusters Benchmark set, we used ESMFold to predict the apo complex structures of each of its proteins. After filtering out 28 complexes for which the corresponding protein structure could not be predicted on an 80GB A100 GPU, we RMSD-aligned the remaining 280 predicted protein structures while optimally weighting each complex’s protein-ligand interface in the alignment. For the DockGen dataset, we refer readers to Appendix G.1.
@@
Sequence identity overlap. Note that for all four of the test datasets described above and listed in Table 1, we do not perform an analysis of the sequence identity overlap between these test datasets’ proteins and those of e.g., the PDBBind (time-split) training dataset, as (1) not all DL-based docking methods use PDBBind as their respective training datasets and (2) leaving the test complexes unfiltered according to sequence identity should, in principle, reflect many real-world use cases of these methods in which several (new) protein targets they are presented with may or may not be similar to what the methods have "seen" during training. Nevertheless, for an investigation of the sequence identity overlap between e.g., the PoseBusters Benchmark set and PDBBind, we refer interested readers to Buttenschoen et al. (2024). Furthermore, in Appendix F, we analyze the different types and frequencies of protein-ligand interactions natively found within the Astex Diverse, PoseBusters Benchmark, DockGen, and CASP15 datasets, respectively, to quantify the diversity of the (predicted) interactions each dataset can be used to evaluate.
@@ 4 Methods and experimental setup
[enumerate] \1 Formats for conventional methods are as follows: \2 Template-based methods such as TULIP are provided with an apo (predicted) protein structure and (fragment) ligand SMILES strings and are tasked with retrieving (PDB template (Bank, 1971)) ligand conformations residing in the same coordinate system as the given (predicted) protein structure following optimal molecular and structural alignment (Hu et al., 2018) with corresponding RDKit conformers of the input (query) ligand SMILES strings, where structural similarity with the query ligands is used to rank-order the selected (PDB template) conformations. \2 Molecular docking tools such as AutoDock Vina, which require specification of protein binding sites, are provided with not only a predicted protein structure but also the centroid coordinates of each (DiffDock-L-)predicted protein-ligand binding site residue. Such binding site residues are classified using a 4 Å protein-ligand heavy atom interaction threshold and using a 25 Å ligand-ligand heavy atom interaction threshold to define a "group" of ligands belonging to the same binding site and therefore residing in the same 25 Å3-sized binding site input voxel for AutoDock Vina. For interested readers, in Appendix G.1, we additionally report results using P2Rank (Krivák and Hoksza, 2018) to predict AutoDock Vina’s binding site centroid inputs.
@@
Formats for generative methods are as follows: \2 DiffDock-L is provided with a predicted protein structure and (fragment) ligand SMILES strings and is then tasked with producing (multiple rank-ordered) ligand conformations (for each fragment) for the given protein. Note that DiffDock-L does not natively support multi-ligand SMILES string inputs, so in this work, we propose a modified inference procedure for DiffDock-L which autoregressively presents each (fragment) ligand SMILES string to the model while providing the same predicted protein structure to the model in each inference iteration (reporting for each complex the average confidence score over all iterations). Notably, as an inference-time modification, this sampling formulation permits multi-ligand sampling yet cannot model multi-ligand interactions directly and therefore often produces ligand-ligand steric clashes. \2 As a single-ligand generative docking method, DynamicBind adopts the same input and output formats as DiffDock-L with the following exceptions: (1) the predicted input protein structure can be modified in response to (fragment) ligand docking; (2) the autoregressive inference procedure we adapted from that of DiffDock-L now provides DynamicBind with its own most recently generated protein structure in each (fragment) ligand inference iteration, thereby providing the model with partial multi-ligand interaction context; and (3) iteration-averaged confidence scores and predicted affinities are reported for each complex. Nonetheless, for both DiffDock-L and DynamicBind, such modified inference procedures highlight the importance in future work of retraining such generative methods directly on multi-ligand complexes to address such inference-time compromises. \2 Lastly, as a natively multi-ligand structure generation model pretrained using various 3D molecular and protein data sources, NeuralPLexer receives as its inputs a (multi-chain) protein sequence, a predicted protein (template) structure, as well as (fragment) ligand SMILES strings. The method is then tasked with producing (multiple rank-ordered) protein-ligand structure conformations for each input complex, using the method’s average predicted per-ligand heavy atom local Distance Difference Test (lDDT) score (Mariani et al., 2013) for rank-ordering.
@@
Multi-ligand evaluation. Following CASP15’s official scoring procedure for protein-ligand complexes (Robin et al., 2023), for multi-ligand targets, we report each method’s percentage of "correct" (binding site-superimposed) ligand conformations (RMSD $\leq$ 2 Å) as well as violin plots of the RMSD and PLI-specific lDDT scores of its protein-ligand conformations across all (fragment) ligands within the benchmark’s multi-ligand complexes (see Appendix G for these plots). Notably, this final metric, referred to lDDT-PLI, allows one to evaluate specifically how well each method can model protein-ligand structural interfaces. We refer readings to Appendix D for formal definitions of these metrics. In the remainder of this work, we will discuss our benchmark’s results and their implications for the development of future complex structure generation methods.
@@ 5 Results and discussions
In this section, we present PoseBench’s results for single and multi-ligand protein-ligand structure generation and discuss their implications for future work. Note that across all the experiments, for generative methods (or methods that use generative inputs to make their predictions), we report their performance metrics in terms of the mean and standard deviation across three independent runs of the method to gain insights into its inter-run stability and consistency. For interested readers, in Appendix C, we report the average runtime and memory usage of each baseline method to determine which methods are the most practical for real-world docking applications.
@@ 5.1 Generalization to new binding pockets implies single-ligand docking performance
We begin our investigations by evaluating the performance of each baseline method for single-ligand docking using the Astex Diverse and PoseBusters Benchmark datasets. Notably, for results on the PoseBusters Benchmark dataset, we perform an additional analysis where we apply post-prediction (fixed-protein) relaxation to each method’s generated ligand conformations using molecular dynamics simulations (Eastman and Pande, 2010), as originally proposed by Buttenschoen et al. (2024). Additionally, for interested readers, in Appendix G.1 we include benchmark results for flexible-protein relaxation as implemented by Lu et al. (2024). As shown in Figures 2 and 3, DiffDock-L achieves the best overall performance across the two datasets both with and without applying relaxation to its generated structures. Closely behind in performance for the PoseBusters Benchmark dataset are DynamicBind and RoseTTAFold-All-Atom following structural relaxation. Interestingly, without relaxation, AutoDock Vina combined with DiffDock-L’s predicted binding pockets achieves the second-best performance on the PoseBusters Benchmark dataset, which suggests that DiffDock-L is currently the only single-ligand deep learning method that presents a better intrinsic understanding of biomolecular physics for docking than conventional modeling tools. For interested readers, in Appendix G.2, we report e.g., pocket-only PoseBusters Benchmark experiments and RMSD violin plots for both the Astex Diverse and PoseBusters Benchmark datasets.
@@ 5.2 Molecule pretraining implies multi-ligand docking performance
Another interesting observation is that TULIP (following structural relaxation) outperforms the docking success rates of single-ligand deep learning docking methods such as DiffDock-L and DynamicBind in the context of multi-ligand docking, which suggests room for future improvement in the multi-ligand modeling capabilities of these recent deep learning baselines. To further inspect each method’s understanding of biomolecular physics for docking, in Figure 5 we report each method’s percentage of predicted complexes (whether correct or not) for which all ligand conformations in the complex are jointly considered valid according to the PoseBusters software suite (i.e., PB-Valid). In short, in the context of multi-ligands, we find that AutoDock Vina followed by DiffDock-L are tied in terms of their PoseBusters validity rates following structural relaxation, with our ensembling consensus baseline (i.e., Ensemble (Con)) as well as DynamicBind shortly behind. As an addendum, we note that NeuralPLexer’s (DynamicBind’s) predictions seem to be frequently selected by Ensemble (Con) for single (multi)-ligand targets, which suggests that NeuralPLexer (DynamicBind) produces the most consistent (i.e., similar) ligand poses for a given single (multi)-ligand protein complex. For interested readers, in Appendix G.3, we report additional results e.g., in terms of lDDT-PLI and RMSD violin plots for both the total available CASP15 targets as well as the publicly available ones.
@@ Conclusions
In this work, we introduced PoseBench, the first deep learning benchmark for practical protein-ligand docking. Experimental results with PoseBench suggest the importance of developing new multi-ligand structure generation methods for enhanced generalization in future work. Moreover, based on these benchmark results, we posit that advances in protein-ligand docking will also likely be driven by advances in modeling macromolecular structures (Abramson et al., 2024) (e.g., by training models on full protein-nucleic acid-ligand complexes), in contrast to current methods that are trained primarily on one type of biomolecular complex (e.g., protein-ligand complexes). Key limitations of this study include its reliance on the accuracy of its predicted protein structures, its limited number of multi-ligand prediction targets available for benchmarking, and its inclusion of only a subset of all available protein-ligand docking baselines to focus on the most recent deep learning algorithms designed specifically for docking. In future work, we aim to expand not only the number of baseline methods but also the number of available multi-ligand targets while maintaining a diverse composition of heterogeneous (ionic) complexes. As a publicly available resource, PoseBench is flexible to accommodate new datasets and methods for protein-ligand structure generation.
@@ Appendix C Compute resources
As a more formal investigation of the computational resources required to run each baseline method in this work, in Table 2 we list the average runtime (in seconds) and peak CPU (GPU) memory usage (in GB) consumed by each method when running them on a 25% subset of the Astex Diverse dataset.
Appendix D Metrics
Table 2: The average runtime (in seconds) and peak memory usage (in GB) of each baseline method on a 25% subset of the Astex Diverse dataset (using an NVIDIA 80GB A100 GPU for benchmarking). The symbol - denotes a result that could not be estimated. Where applicable, an integer enclosed in parentheses indicates the number of samples drawn from a particular (generative) baseline method.
Method Runtime (s) CPU Memory Usage (GB) GPU Memory Usage (GB)
DiffDock-L (40) 130.53 9.67 63.07
FABind 4.01 5.00 8.44
DynamicBind (40) 187.00 5.36 79.11
NeuralPLexer (40) 223.65 11.31 42.61
RoseTTAFold-All-Atom 862.60 49.78 78.97
TULIP - - -
DiffDock-L-Vina 13.05 0.80 0.00
P2Rank-Vina 17.83 2.13 0.00
Ensemble (Con) - - -
In this work, we reference two key metrics in the field of structural bioinformatics: RMSD and lDDT. The RMSD between a predicted ligand 3D conformation (with atomic positions $\hat{x}_{i}$ for each of the ligand’s $n$ heavy atoms) and the ground-truth conformation ($x_{i}$) is defined as:
$\text{RMSD}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\lVert\hat{x}_{i}-x_{i}\rVert^{2}}.$ (1)
The lDDT score, which is commonly used to compare predicted and ground-truth protein 3D structures, is defined as:
$\text{lDDT}=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{4}\sum_{k=1}^{4}\left(\frac{1}{|\mathcal{N}_{i}|}\sum_{j\in\mathcal{N}_{i}}\Theta(|\hat{d}_{ij}-d_{ij}|<\Delta_{k})\right),$ (2)
where $N$ is the total number of heavy atoms in the ground-truth structure; $\mathcal{N}_{i}$ is the set of neighboring atoms of atom $i$ within the inclusion radius $R_{o}=15\ \AA$ in the ground-truth structure, excluding atoms from the same residue; $\hat{d}_{ij}$ ($d_{ij}$) is the distance between atoms $i$ and $j$ in the predicted (ground-truth) structure; $\Delta_{k}$ are the distance tolerance thresholds (i.e., 0.5 Å, 1 Å, 2 Å, and 4 Å); $\Theta(x)$ is a step function that equals 1 if x is true, and 0 otherwise; and $|\mathcal{N}_{i}|$ is the number of neighboring atoms for atom $i$.
As originally proposed by Robin et al. [2023], in this study, we adopt the PLI-specific variant of lDDT, which calculates lDDT scores to compare predicted and ground-truth protein-ligand complex structures following optimal structural alignment of the predicted and ground-truth protein-ligand binding pockets.
Appendix E Documentation for datasets
@@
E.1 Astex Diverse Set - Single-Ligand Docking (Difficulty: Easy)
@@
E.2 PoseBusters Benchmark Set - Single-Ligand Docking (Difficulty: Intermediate)
@@
E.3 DockGen Set - Single-Ligand Docking (Difficulty: Challenging)
@@
E.4 CASP15 Set - Multi-Ligand Docking (Difficulty: Challenging)
@@
Appendix F Analysis of protein-ligand interactions
@@
Appendix G Additional results
@@ G.1 DockGen results
Benchmark results. Figures 8 and 9 show that DiffDock-L, NeuralPLexer, and Ensemble (Con) provide the best pocket generalization capabilities compared to all other baseline methods. In particular, Ensemble (Con) seems to provide the most stable performance in this setting. Moreover, the results for DiffDock-L-Vina and P2Rank-Vina suggest that DiffDock-L generalizes to new binding pockets slightly better than P2Rank for conventional docking with AutoDock-Vina. Additionally, DiffDock-L’s results with protein-flexible relaxation applied (i.e., DiffDock-L (Relax-P)) demonstrate that fixed-protein relaxation (albeit unideal from a theoretical e.g., protein side chain perspective [Wankowicz et al., 2022]) yields less accuracy degradation to DiffDock-L’s original ligand conformations compared to side chain-flexible relaxation. However, we note that none of the baseline methods generate any PB-valid ligand conformations, suggesting that such "correct" poses are approximately accurate yet physically implausible in certain measurable ways.
@@
G.2 Expanded Astex & PoseBusters results
G.2.1 Pocket-only PoseBusters results